# From User Surveys to Telemetry-Driven Agents: Exploring the Potential of Personalized Productivity Solutions

### v1

@@ 4.4. AI Productivity Agent Performance
In this section, we delve into the performance and capabilities of the AI productivity agent, examining the aspects of its interaction with users. We analyze the perceived capabilities of the agent, chat metrics, and productivity-related themes that emerged from users’ questions. Our goal is to gain insights into the effectiveness of the AI productivity agent.
@@
Table 3. Perceived capabilities of the interactive agent. The responses are categorized as follows: SD for Strongly Disagree, SWD for Somewhat Disagree, NAND for Neither Agree Nor Disagree, SWA for Somewhat Agree, SA for Strongly Agree, and N/A for Non-applicable. We bold the group that has the majority of the responses – either Agree (SA, WA) or Disagree (SD, SWD) or Neutral (NAND, N/A).
@@
The agent provided useful tips and strategies for improving productivity
1 (2.5%) 1 (2.5%) 6 (15%) 15 (37.5%) 17 (42.5%) 0 (0%)
The agent was able to understand my productivity challenges
0 (0%) 1 (2.5%) 8 (20%) 19 (47.5%) 9 (22.5%) 3 (7.5%)
@@
The agent’s empathy and tone helped me feel more comfortable discussing my productivity concerns
2 (5%) 5 (12.5%) 10 (25%) 9 (22.5%) 12 (30%) 2 (5%)
The agent personalized its suggestions based on my situation and needs
1 (2.5%) 2 (5%) 6 (15%) 16 (40%) 14 (35%) 1 (2.5%)
The agent provided practical solutions to overcome productivity obstacles
3 (7.5%) 1 (2.5%) 5 (12.5%) 16 (40%) 13 (32.5%) 2 (5%)
The agent used a professional and direct language and tone
1 (2.5%) 0 (0%) 0 (0%) 13 (32.5%) 26 (65%) 0 (0%)
The agent proactively suggested ways to improve my productivity
2 (5%) 0 (0%) 5 (12.5%) 17 (42.5%) 15 (37.5%) 1 (2.5%)
The agent was annoying, aggressive or pushy
26 (65%) 7 (17.5%) 3 (7.5%) 1 (2.5%) 2 (5%) 1 (2.5%)
The agent was being too intrusive
22 (55%) 11 (27.5%) 4 (10%) 2 (5%) 0 (0%) 1 (2.5%)
The agent was overly critical
20 (50%) 11 (27.5%) 5 (12.5%) 1 (2.5%) 0 (0%) 3 (7.5%)
@@
The agent seems to be able to easily work out what I might want to talk about
1 (2.5%) 1 (2.5%) 10 (25%) 17 (42.5%) 10 (25%) 1 (2.5%)
@@
The agent seems to have a difficult time seeing things from my point of view
8 (20%) 11 (27.5%) 12 (30%) 4 (10%) 3 (7.5%) 2 (5%)
@@
The agent seems to vary its conversational style to accommodate my mood or disposition
1 (2.5%) 3 (7.5%) 17 (42.5%) 7 (17.5%) 1 (2.5%) 11 (27.5%)
The agent seems to follow what I say and accurately reflects its understanding to me
1 (2.5%) 2 (5%) 3 (7.5%) 17 (42.5%) 17 (42.5%) 0 (0%)
@@
The agent demonstrated the traits of being knowledgeable, trustworthy, transparent, and responsive
1 (2.5%) 3 (7.5%) 5 (12.5%) 17 (42.5%) 14 (35%) 0 (0%)
@@
The agent’s use of telemetry data enhanced its ability to provide personalized and empathetic support
1 (2.5%) 5 (12.5%) 3 (7.5%) 17 (42.5%) 12 (30%) 2 (5%)
@@
29 participants (72.5%) agreed that the use of telemetry data in the agent enhanced its ability to provide personalized and empathetic support. To design the agent as knowledgeable, trustworthy, and transparent based on the user survey, 31 participants (77.5%) felt that the agent demonstrated these traits. Additionally, 34 participants (84%) agreed that the agent followed directions well. However, only 8 participants (20%) found the agent adaptive in its conversation style, with 17 (42.5%) remaining neutral and 11 (27.5%) considering it non-applicable. The agent performed well in avoiding negative traits identified in the user survey (we instructed the agent through its system prompt to avoid such negative traits), as 31 participants (77.5%) disagreed that it was overly critical, 33 (82.5%) disagreed that it was intrusive, and 33 (82.5%) disagreed that it was annoying, aggressive, or pushy. Nevertheless, the agent’s ability to see things from the user’s point of view received mixed responses, with only 19 participants (47.5%) disagreeing, 7 (17.5%) agreeing, and 12 (30%) remaining neutral. Other findings indicated that the agent proactively suggested productivity improvements (80%), provided practical solutions to overcome productivity barriers (72.5%), and offered useful tips and strategies (82.5%). Participants also felt that the agent personalized its suggestions based on their situation and needs (75%) and understood their productivity challenges (70%).
@@ 6. Discussion
Our study adopted a comprehensive two-phase approach to delve into user preferences for productivity agents and devise personalized solutions tailored to these preferences. The initial phase entailed a user survey to pinpoint the ideal persona for a productivity agent, examining individuals’ expectations and preferences regarding its features and behaviors. The subsequent phase centered on crafting an AI-based productivity agent, assessing its performance relative to alternative modes of telemetry data presentation, such as dashboard visualization and personalized narratives.
@@
The user survey in phase one investigated various aspects of productivity, communication style, agent approach, personality traits, personalization, and privacy. The findings indicated that progress on scheduled tasks, resolving unexpected tasks, and assisting colleagues significantly contributed to participants’ sense of productivity. Most participants believed a productivity agent would be advantageous and that its personality would moderately or significantly influence its effectiveness. The participants desired AI agent assistance in areas like meeting summaries, documentation writing, start-of-day briefs, work summaries, and information retrieval. These results are in line with  Kimani et al. (2019), who discovered that emails and meetings were the primary distractions affecting information workers’ productivity. Both studies found that progressing or completing scheduled tasks was the top measure of productivity for participants. However, an intriguing difference emerged in the areas where respondents in the Kimani et al. study sought assistance. They primarily indicated a preference for help with reminders, task scheduling, breaking down larger tasks into manageable chunks, and avoiding distractions. While several of these aspects appear on our list of desired support areas, they are not ranked as highly as in the Kimani et al. study. This discrepancy raises the question of whether the areas where people need assistance from AI productivity agents vary depending on their understanding of the agent’s capabilities. This notion brings up an interesting point that the responses in our study could have been shaped more by participants’ familiarity with the capabilities of LLMs, as portrayed by media and social discourse, rather than by their direct experiences with these technologies. There is a possibility that the hype and discussions about LLM-powered tools in public and online forums might have swayed participants’ perceptions, aligning them with tasks where AI is commonly viewed as beneficial. Our research did not explicitly separate preferences influenced by hands-on LLM use from those impacted by societal narratives. However, acknowledging this influence is critical, as it implies that user expectations for AI support can be shaped not only by personal use but also by the prevailing stories about AI’s potential in the wider world. This insight sheds more light on user attitudes towards AI productivity tools and their evolving expectations. It is important to note that the discrepancy in the areas where the respondents sought assistance could also stem from simply demographic variations in the research subjects. Kimani et al.’s study focused on information workers in positions such as program managers, administrative assistants, researchers, developers, engineers, and broadcast producers. Given that our participants report similar job roles, it’s unlikely that occupational differences are a major factor in the discrepancy but other demographic factors could have played a role.
In terms of communication style, participants preferred a communication tone that balanced casual and formal styles, with direct and concise messages. This preference might be attributed to the modern work environment’s emphasis on balancing professionalism and fostering a friendly atmosphere, which can promote collaboration and better working relationships. The preference for direct and to-the-point messages might also be a result of the fast-paced nature of information work, where employees often have limited time to spend on interactions and prefer clear, concise information. The majority favored a proactive agent that displayed traits like being knowledgeable, trustworthy, and transparent, and took on the role of an assistant or expert rather than a supervisor. Regarding personalization and privacy, most participants preferred moderate to high personalization in the productivity agent and were comfortable sharing high-level telemetry data for better guidance. However, some participants expressed concerns about sharing personal information. These findings from the first phase provided valuable insights for designing a personalized productivity agent that effectively addresses users’ preferences and requirements.
@@
Interestingly, approximately 23% of participants, as detailed in Table 8 (Appendix H), expressed neutrality or did not believe productivity assistants would be beneficial for them. This subset highlighted several key concerns. One of the participants, for instance, differentiated between overall productivity and specific motivational challenges, noting that while they generally did not struggle with productivity, they occasionally faced difficulties in motivation, particularly with tasks that were unengaging or poorly defined. This gives us the impression that the participant believes such motivational challenges to be distinct from their overall productivity and not something they expect a productivity assistant to effectively address. Additionally, there were reservations about the factual accuracy of AI responses, especially for fact-based questions, reflecting a broader concern about the reliability of AI tools. Further, a notable apprehension was the potential over-reliance on AI and its impact on personal growth and skill development. Participants expressed uncertainty about their ability to challenge themselves if they became too dependent on AI assistance. Moreover, skepticism about the practical usefulness of AI in conversations was evident, with some participants emphasizing the need for real data to validate the effectiveness of such tools. These insights suggest that while AI agents can offer structural support, they might be less equipped to address intrinsic motivational factors and that there is a need for continual improvement in AI reliability and user empowerment. Understanding these perspectives is essential for refining the design of AI productivity tools and enhancing their adoption and impact in the workplace.
In the second phase of the study, we developed the AI productivity agent based on the insights obtained from the initial user survey. We then compared the AI agent’s performance with two alternative methods of presenting telemetry data: a visual dashboard and an AI-based personalized narrative. User testing allowed for a comprehensive assessment of the AI productivity agent and the alternatives, considering factors such as task assistance, personalization, and overall user experience. The interaction data highlights the difference in user engagement and interaction patterns across the various presentation modes, indicating that users spent more time interacting with the AI productivity agent compared to the dashboard and personalized narrative. The AI productivity agent also offered more comprehensive responses, with a higher word count per message. The notable engagement duration with the AI agent, the longest among the tools evaluated, presents an interesting dimension to our findings. This prolonged interaction could be partly due to the AI agent’s capacity to provide more comprehensive and detailed responses, which inherently requires more time for users to read and digest. Such in-depth responses suggest a depth of interaction and a level of detail that may not be present in other tools, potentially leading to a richer user experience. However, it’s also important to recognize that some of this extended engagement might be attributed to the response generation time inherent in AI systems. Delays in producing responses, combined with the detailed nature of these responses, could have contributed to the longer interaction times. While longer engagement times do not unilaterally equate to higher productivity or satisfaction, they do suggest a level of user interest and investment in the tool. This engagement could also reflect the users’ exploration of the AI agent’s capabilities, a positive sign of their willingness to integrate advanced technology into their productivity workflows. Future research should aim to untangle these factors – response time and response depth – to better understand their individual and combined impacts on user experience, engagement, and productivity.
The AI agent demonstrated promise in providing specific, tailored insights and recommendations. Most of the participants agreed that the use of telemetry data in the agent enhanced its ability to provide personalized and empathetic support, and they felt that the agent demonstrated desired traits such as knowledgeability, trustworthiness, and transparency. It is important to note that in addition to customizing responses, the availability of telemetry data can provide individuals with valuable insights into their work patterns and habits, fostering self-reflection. Previous research has demonstrated that the presence of such data can greatly enhance one’s ability to engage in self-reflection (McDuff et al., 2012; Kocielnik et al., 2018). The agent also performed well in avoiding negative traits identified in the user survey, such as being overly critical, intrusive, and annoying. However, only a minority of participants found the agent to be adaptive in its conversation style. This could be because the study was not longitudinal, and participants only interacted with the agent for a few minutes. In such a limited time, they might not have had the ability to decide about the adaptiveness of the conversation style. Future studies could involve an extended interaction period, allowing users to better understand and evaluate the agent’s adaptiveness.
Participants also expressed a preference for a combination of the visual appeal and ease of understanding offered by the dashboard and the in-depth, personalized insights provided by the interactive chat agent. This preference suggests that users desire a comprehensive productivity tool that combines the strengths of both methods. One possible reason for this preference is that visual representations of data can facilitate the easy comprehension and identification of trends, while the interactive chat agent can provide more detailed information and recommendations tailored to the user’s specific situation. There has been some prior research suggesting that visualizations can be better for browsing and getting oversights, while text summaries are better for targeted search for information (Smith et al., 2006). To further improve the interactive chat agent and better cater to user preferences, future work should focus on integrating the text and visualizations to address both needs, in addition to benchmarking, and other enhancements. Incorporating visual elements into the chat agent’s responses can help users more easily understand and interpret the data, catering to the preference for a combination of dashboard visualization and interactive chat agent. Benchmarking functionality, allowing users to compare their productivity data against their team, organization, or company averages, can provide valuable context and motivation for improvement. The personalized narrative method received mixed reviews, possibly due to the varying preferences of users when it comes to data presentation. Some may prefer a visually engaging presentation, while others might appreciate the succinct and informative nature of the narrative. This highlights the importance of offering multiple presentation methods or a hybrid approach to accommodate diverse user preferences. In summary, the study findings suggest that the interactive chat agent shows promise in providing tailored insights and recommendations, but there is room for improvement to enhance its effectiveness and user experience. For instance, in the example provided in Appendix  I Figure 8(b), there appear to be contradictions in the AI agent’s responses regarding email management strategies. Point 2 suggests specific slots for emails whereas Point 4 suggests spreading emails more evenly throughout the day. Such inconsistencies might have impacted participants’ perceptions and ratings of the agent, and highlight the need for further refinement in the AI’s response generation process. Improving the prompts and the underlying logic of the AI could help ensure more coherent and consistent advice, potentially leading to better user satisfaction and trust in the AI agent. Future iterations of the tool will focus on enhancing the clarity and consistency of the AI’s recommendations. In addition, based on our findings, it seems that to create a more effective and user-centric productivity-enhancing tool, incorporating design principles like semantic zooming and progressive disclosure can be beneficial. Semantic zooming  (Shneiderman, [n. d.]) allows users to view data at varying levels of detail based on their zoom level, presenting both visually appealing overviews and in-depth insights as needed. This could help users better understand relationships between productivity metrics and identify patterns or trends, promoting informed decision-making and productivity improvements. Progressive disclosure (Nielsen, 2023) presents information in a layered manner, initially offering essential information and revealing more details as needed. This approach caters to preferences for direct and concise messages, allowing users to control the level of detail they receive and ensuring the agent’s guidance remains relevant and useful. Integrating these design principles into the AI productivity agent may improve its effectiveness and user experience, striking a balance between visually engaging overviews and in-depth, personalized insights. Future research should consider the benefits of incorporating these principles into productivity agent development and explore additional techniques for creating effective and user-centric productivity-enhancing tools.
Our study also explored the capabilities of GPT-4 in enhancing the customization and delivery of productivity data from platforms like Viva Insights through an AI chatbot/agent. Its advanced natural language processing capabilities allowed us to transform standard productivity metrics into more personalized, interactive, and accessible advice. By leveraging the sophisticated context understanding of GPT-4, we were able to provide a conversational interface that offered more than just data visualization; it provided tailored recommendations based on individual user patterns and preferences. This approach elevates the functionality of traditional productivity tools like Viva Insights, making the insights more actionable for the user. However, our study also recognized the importance of balancing personalization with user privacy. In designing our productivity agent, we adhered to privacy-by-design principles, using only high-level telemetry data and ensuring user consent for data usage. GPT-4’s efficiency in generating relevant and personalized responses without delving into sensitive personal details was key to maintaining this balance. This approach highlights the potential of LLMs to enhance user experience while respecting privacy concerns. Our study thus contributes to the growing body of research on AI-enhanced productivity tools, demonstrating the potential of advanced language models like GPT-4 in transforming how productivity insights are delivered and interacted with. Future research should aim to further isolate and examine the effects of presentation and functionality variations to provide clearer insights into the most effective ways of enhancing user productivity.
7. Limitations

### v2

@@ 4.4. AI Productivity Agent Performance
In this section, we look into the performance and capabilities of the AI productivity agent, examining the aspects of its interaction with users. We analyze the perceived capabilities of the agent, chat metrics, and productivity-related themes that emerged from users’ questions. Our goal is to gain insights into the effectiveness of the AI productivity agent.
@@
Table 3. Perceived capabilities of the interactive agent. The responses are categorized as follows: SD for Strongly Disagree, SWD for Somewhat Disagree, NAND for Neither Agree Nor Disagree, SWA for Somewhat Agree, SA for Strongly Agree, and N/A for Non-applicable. We bold the group that has the majority of the responses – either Agree (SA, SWA) or Disagree (SD, SWD) or Neutral (NAND, N/A).
@@
The agent provided useful tips and strategies for improving productivity 1 (2.5%) 1 (2.5%) 6 (15%) 15 (37.5%) 17 (42.5%) 0 (0%)
The agent was able to understand my productivity challenges 0 (0%) 1 (2.5%) 8 (20%) 19 (47.5%) 9 (22.5%) 3 (7.5%)
@@
The agent’s empathy and tone helped me feel more comfortable discussing my productivity concerns 2 (5%) 5 (12.5%) 10 (25%) 9 (22.5%) 12 (30%) 2 (5%)
The agent personalized its suggestions based on my situation and needs 1 (2.5%) 2 (5%) 6 (15%) 16 (40%) 14 (35%) 1 (2.5%)
The agent provided practical solutions to overcome productivity obstacles 3 (7.5%) 1 (2.5%) 5 (12.5%) 16 (40%) 13 (32.5%) 2 (5%)
The agent used a professional and direct language and tone 1 (2.5%) 0 (0%) 0 (0%) 13 (32.5%) 26 (65%) 0 (0%)
The agent proactively suggested ways to improve my productivity 2 (5%) 0 (0%) 5 (12.5%) 17 (42.5%) 15 (37.5%) 1 (2.5%)
The agent was annoying, aggressive or pushy 26 (65%) 7 (17.5%) 3 (7.5%) 1 (2.5%) 2 (5%) 1 (2.5%)
The agent was being too intrusive 22 (55%) 11 (27.5%) 4 (10%) 2 (5%) 0 (0%) 1 (2.5%)
The agent was overly critical 20 (50%) 11 (27.5%) 5 (12.5%) 1 (2.5%) 0 (0%) 3 (7.5%)
@@
The agent seems to be able to easily work out what I might want to talk about 1 (2.5%) 1 (2.5%) 10 (25%) 17 (42.5%) 10 (25%) 1 (2.5%)
@@
The agent seems to have a difficult time seeing things from my point of view 8 (20%) 11 (27.5%) 12 (30%) 4 (10%) 3 (7.5%) 2 (5%)
@@
The agent seems to vary its conversational style to accommodate my mood or disposition 1 (2.5%) 3 (7.5%) 17 (42.5%) 7 (17.5%) 1 (2.5%) 11 (27.5%)
The agent seems to follow what I say and accurately reflects its understanding to me 1 (2.5%) 2 (5%) 3 (7.5%) 17 (42.5%) 17 (42.5%) 0 (0%)
@@
The agent demonstrated the traits of being knowledgeable, trustworthy, transparent, and responsive 1 (2.5%) 3 (7.5%) 5 (12.5%) 17 (42.5%) 14 (35%) 0 (0%)
@@
The agent’s use of telemetry data enhanced its ability to provide personalized and empathetic support 1 (2.5%) 5 (12.5%) 3 (7.5%) 17 (42.5%) 12 (30%) 2 (5%)
@@
29 participants (72.5%) agreed that the use of telemetry data in the agent enhanced its ability to provide personalized and empathetic support. To design the agent as knowledgeable, trustworthy, and transparent based on the user survey, 31 participants (77.5%) felt that the agent demonstrated these traits. Additionally, 34 participants (84%) agreed that the agent followed directions well. However, only 8 participants (20%) found the agent adaptive in its conversation style to accommodate participants’ mood, with 17 (42.5%) remaining neutral and 11 (27.5%) considering it non-applicable. We performed a Wilcoxon Signed-Rank Test ($\alpha$ = 0.10) to determine if participants significantly agreed or disagreed compared to a neutral response. We found significant differences from neutral responses in all questions except this one i.e., change in conversation style to accommodate participants’ mood or disposition (p-value > 0.10). The agent performed well in avoiding negative traits identified in the user survey (we instructed the agent through its system prompt to avoid such negative traits), as 31 participants (77.5%) disagreed that it was overly critical, 33 (82.5%) disagreed that it was intrusive, and 33 (82.5%) disagreed that it was annoying, aggressive, or pushy. Nevertheless, the agent’s ability to see things from the user’s point of view received mixed responses, with only 19 participants (47.5%) disagreeing, 7 (17.5%) agreeing, and 12 (30%) remaining neutral. Other findings indicated that the agent proactively suggested productivity improvements (80%), provided practical solutions to overcome productivity barriers (72.5%), and offered useful tips and strategies (82.5%). Participants also felt that the agent personalized its suggestions based on their situation and needs (75%) and understood their productivity challenges (70%).
@@ 6. Discussion
Our study employed a two-phase approach to explore productivity agent design: first investigating user preferences through surveys to define an ideal agent persona, then developing and evaluating an AI-based productivity agent against alternative telemetry data presentation modes like dashboards and narratives. The user survey in phase one examined aspects of productivity, communication style, agent traits, personalization, and privacy. Findings showed that participants valued progress on scheduled tasks, resolving unexpected tasks, and assisting colleagues as key productivity indicators. Most participants supported the idea of a productivity agent and believed its personality would significantly impact effectiveness. They sought AI assistance primarily for meeting summaries, documentation, start-of-day briefs, work summaries, and information retrieval. These results partially align with  Kimani et al. (2019), who found emails and meetings were primary productivity distractions, and completing scheduled tasks was the top productivity measure. However, an interesting difference emerged: Kimani et al.’s participants prioritized help with reminders, task scheduling, and distraction management, while our participants ranked these lower. This discrepancy might reflect evolving user expectations shaped by media coverage and public discourse about LLM capabilities rather than direct experience. There is a possibility that the hype and discussions about LLM-powered tools in public and online forums might have swayed participants’ perceptions, aligning them with tasks where AI is commonly viewed as beneficial. While our study did not explicitly distinguish between preferences influenced by hands-on LLM use versus societal narratives, this potential influence suggests that user expectations for AI support are shaped by both personal experience and broader technological narratives. Although both studies involved similar job roles (program managers, developers, engineers), other demographic factors might explain these different preferences. Regarding agent characteristics, participants preferred a balance of casual and formal communication with direct, concise messages—reflecting modern workplace needs for efficient yet relationship-building communication. They favored knowledgeable, transparent assistants over supervisory roles, and most were comfortable sharing basic data for personalization despite some privacy concerns. These preferences informed our productivity agent design in phase two.
@@
Interestingly, approximately 23% of participants, as detailed in Table 8 (Appendix H), expressed neutrality or did not believe productivity assistants would be beneficial for them. Their concerns centered around several key issues. One participant notably distinguished between overall productivity and motivational challenges, suggesting that while they managed productivity well, they struggled with motivation for unengaging or poorly defined tasks—a challenge they believed was beyond the scope of productivity assistants. Participants also raised concerns about AI reliability, particularly regarding factual accuracy, and worried about over-dependence on AI potentially hampering personal growth and skill development. Some expressed skepticism about AI’s practical value in conversations, emphasizing the need for empirical validation of such tools’ effectiveness. These insights suggest that while AI agents might excel at providing structural support, they may fall short in addressing intrinsic motivational factors. Understanding these reservations is crucial for improving AI productivity tools’ design and increasing their workplace adoption (Suh et al., 2023; Das Swain et al., 2024; Saha and Chandrasekharan, 2024).
In the second phase of the study, we developed the AI productivity agent based on insights from our initial survey and compared its performance with two alternative methods of presenting telemetry data: a visual dashboard and an AI-based personalized narrative. The interaction data revealed significant differences in user engagement patterns, with users spending notably more time with the AI productivity agent compared to other presentation modes. The AI agent provided more comprehensive responses, with a higher word count per message. This extended engagement duration presents an interesting dimension to our findings. While it could reflect the agent’s ability to provide detailed, comprehensive responses requiring more time to digest, it might also be partially attributed to inherent AI response generation delays. Though longer engagement times do not necessarily indicate higher productivity or satisfaction, they suggest significant user interest and willingness to explore AI capabilities in productivity workflows. The depth of interaction and detailed responses point to a potentially richer user experience not present in other tools. Future research should investigate the distinct impacts of response time versus response depth on user experience, engagement, and productivity. The AI agent showed promise in providing tailored insights and recommendations, with most participants agreeing that telemetry data enhanced its ability to provide personalized and empathetic support. Users recognized the agent’s demonstration of desired traits like knowledgeability, trustworthiness, and transparency, while also noting its success in avoiding negative characteristics identified in the survey (being overly critical, intrusive, or annoying). Beyond customizing responses, the telemetry data served to foster self-reflection, aligning with previous research showing how such data can enhance reflective practices  (McDuff et al., 2012; Kocielnik et al., 2018). However, only a few participants found the agent adaptive in its conversation style, possibly due to the study’s brief interaction period. Future longitudinal studies could better evaluate the agent’s conversational adaptiveness by allowing extended user interactions.
Participants expressed a preference for combining the dashboard’s visual appeal and ease of understanding with the chat agent’s in-depth, personalized insights. This preference aligns with prior research suggesting that visualizations excel at providing overviews and trend identification, while text summaries better serve targeted information searches (Smith et al., 2006). The complementary nature of these approaches—visual data representation for quick comprehension and interactive dialogue for detailed, personalized recommendations—indicates users’ desire for a comprehensive productivity tool that leverages both methods’ strengths. Future improvements should focus on integrating visualizations into the chat agent’s responses and adding benchmarking functionality to allow users to compare their productivity metrics against team or organizational averages. The mixed reception of the personalized narrative method underscores the importance of supporting diverse user preferences through multiple presentation methods or a hybrid approach. In summary, while the interactive chat agent shows promise in providing tailored insights and recommendations, our analysis revealed areas for improvement. For example, we found inconsistencies in the agent’s advice, as illustrated in Appendix I Figure 9(b), where the agent contradicts itself on email management strategies—suggesting both specific time slots and even distribution throughout the day. Such contradictions may have influenced participants’ perceptions and ratings of the agent, highlighting the need for more refined prompt engineering and improved response generation logic. Future iterations will prioritize enhancing the coherence and consistency of AI recommendations to build better user satisfaction and trust.
Our findings suggest that incorporating design principles like semantic zooming  (Shneiderman, 1996) and progressive disclosure  (Nielsen, 2023) could enhance the effectiveness of productivity tools. Semantic zooming would allow users to explore data at varying levels of detail, facilitating better understanding of relationships between productivity metrics and patterns. Progressive disclosure would present information in layers, starting with essential details and revealing more as needed, aligning with users’ preferences for concise yet detailed guidance. Integrating these principles into future AI productivity agents could help strike an optimal balance between high-level overviews and in-depth insights, while maintaining user control over information density. Our study also explored GPT-4’s capabilities in transforming productivity data from platforms like Viva Insights into more personalized, interactive advice through an AI chatbot. GPT-4’s advanced natural language processing abilities enabled us to create a conversational interface that went beyond simple data visualization, offering tailored recommendations based on individual work patterns. This approach enhanced traditional productivity tools by making insights more actionable for users. While developing the agent, we prioritized user privacy by adhering to privacy-by-design principles, using only high-level telemetry data with explicit user consent. Our research demonstrates how LLMs can enhance user experience while maintaining privacy, contributing to the growing body of work on AI-enhanced productivity tools. Future research should examine different presentation and functionality variations to identify the most effective approaches for enhancing user productivity.
7. Limitations & Future Work
Our study has several limitations worth noting. The sample size and demographic composition may not fully represent the diversity of user preferences and productivity patterns across different professional contexts. The relatively short duration of our study also limited participants’ ability to fully adapt to and evaluate the different presentation modes, potentially affecting the long-term validity of our findings. While our study provides insights into how different personal informatics tools support productivity-related reflection and action, we acknowledge that we did not measure direct productivity improvements through quantitative metrics such as task completion rates, distraction reduction, or comparative efficiency measures. A longitudinal study examining how each presentation mode influences behavior change and productivity outcomes would be valuable for validating the systems’ effectiveness. It is also important to acknowledge how prompt variations can influence GPT-4’s outputs. The model’s sensitivity to prompt phrasing and structure means that even subtle changes could yield significantly different responses, potentially affecting the consistency and comparability of our findings.